OpenStudy (anonymous):

Which of the following systems of inequalities would produce the region indicated on the graph? (will post graph and answer choices in a sec!!!) ***PLease explain!! :) Thank youuuu :)

OpenStudy (anonymous):

GRAPH

OpenStudy (anonymous):

Answer choices A,B,C,D (top to bottom)

jimthompson5910 (jim_thompson5910):

we have 2 lines here

jimthompson5910 (jim_thompson5910):

we can find the equations of these lines by locating the points on the lines

OpenStudy (anonymous):

mhmm and one is solid and one is dotted... do we need to know that part??

jimthompson5910 (jim_thompson5910):

so for instance, the line sloping downward has the points (0,4) and (4,0)

jimthompson5910 (jim_thompson5910):

yes that's important info as well

OpenStudy (anonymous):

yup i see that :)

jimthompson5910 (jim_thompson5910):

what is the equation of the line that goes through (0,4) and (4,0)

OpenStudy (anonymous):

it goes thru (0,2) ??

jimthompson5910 (jim_thompson5910):

that's a different line

OpenStudy (anonymous):

ohhh okay nvm hahaa hold on lemme see

OpenStudy (anonymous):

slope is -1 right?

jimthompson5910 (jim_thompson5910):

yes, slope is -1

jimthompson5910 (jim_thompson5910):

for the solid line

OpenStudy (anonymous):

yup :) and is the y intercept 4?? the y-int is off the vertical axis (y axis) right??

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

so the equation is y = ??x + ??

OpenStudy (anonymous):

so the equation for the solid line is y=-x+4 (i don't need to write -1x right? -x is fine?)

jimthompson5910 (jim_thompson5910):

-x is just fine

jimthompson5910 (jim_thompson5910):

ys, y = -x+4 is the equation of the boundary line

jimthompson5910 (jim_thompson5910):

yes*

jimthompson5910 (jim_thompson5910):

so because the shading is below this solid boundary line, you use the \(\Large \le\) symbol

jimthompson5910 (jim_thompson5910):

to get \(\Large y \le -x+4\)

OpenStudy (anonymous):

haha typos... we hate them! lol :) oh okay i see :) that makes sense :) what next?

jimthompson5910 (jim_thompson5910):

ys i hate typos lol (totally intentional...hmm probably killed the joke though)

jimthompson5910 (jim_thompson5910):

use the same idea to find the inequality for the second piece

OpenStudy (anonymous):

lawlz i see what you did there! hahaa :) good one :) and so we are at the second line now?? the dotted one?

jimthompson5910 (jim_thompson5910):

yes, what is the equation of that boundary line?

OpenStudy (anonymous):

the dotted one?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

ummm so the point is (0,2) right?? so the y-intercept is 2 yes?

jimthompson5910 (jim_thompson5910):

good so far, what's another point? or what's the slope?

OpenStudy (anonymous):

and the slope looks like it is 1 right? is the slop=1?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation is y = x+2

OpenStudy (anonymous):

another point would be (1,3)

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

yes, another point is (1,3)

jimthompson5910 (jim_thompson5910):

the line is dashed and we're shading below it, so the inequality is y < x+2

jimthompson5910 (jim_thompson5910):

hmm something seems weird...it's like some of your messages aren't getting through?

OpenStudy (anonymous):

so is this the answer?? answer A?? \[y \le -x +4;y <x +2;y \ge0,x \ge0\]

OpenStudy (anonymous):

is that right?? answer A?

jimthompson5910 (jim_thompson5910):

exactly, A is the answer

OpenStudy (anonymous):

and yeah it's weird right?? i had to retype that :( oooh!! yay :) thanks!!! :)

jimthompson5910 (jim_thompson5910):

lame, i hate when that happens

OpenStudy (anonymous):